OpenStudy (anonymous):

helpp. Solve the proportion. 10/a = 15/30

OpenStudy (anonymous):

Cross multiply: 10*30 = 15*a 300 = 15a 300/15 = a ? = a

OpenStudy (anonymous):

a=20?

OpenStudy (anonymous):

Exactly. You're awesome! <3 :D

OpenStudy (anonymous):

Thankyou !!!